# Massive  Index  Modulation  with  Combinatorics-Free Detection  for  Integrated  Sensing  and  Communications

### v1

@@ Abstract
Integrated sensing and communications (ISAC) and index modulation (IM) are promising technologies for beyond fifth generation (B5G) and sixth generation (6G) systems. While ISAC enables new applications, IM is attractive for its inherent energy and spectral efficiencies. In this article we propose massive IM as an enabler of ISAC, by considering transmit signals with information conveyed through the indexation of the resources utilized in their transmission, and pilot symbols exploited for sensing. In order to overcome the complexity hurdle arising from the large sizes of IM codebooks, we propose a novel message passing (MP) decoder designed under the Gaussian belief propagation (GaBP) framework exploiting a novel unit vector decomposition (UVD) of IM signals with purpose-derived novel probability distributions. The proposed method enjoys a low decoding complexity that is independent of combinatorial factors, while still approaching the performance of unfeasible state-of-the-art (SotA) search-based methods. The effectiveness of the proposed approach is demonstrated via complexity analysis and numerical results for piloted generalized quadrature spatial modulation (GQSM) systems of large sizes (up to 96 antennas).
@@
Integrated sensing and communications, index modulation, quadrature spatial modulation, massive multiple-input multiple-output, message passing, belief propagation, unit vector decomposition, beyond fifth generation wireless systems.
@@ I Introduction
Integrated sensing and communications (ISAC) has recently been recognized as a new key technology for beyond fifth generation (B5G) and sixth generation (6G) wireless communications systems [3, 4, 5, 6]. But although ISAC is expected to enable new applications, the fundamental trade-off between sensing and communication functionalities [7] is also certain to put pressure on the equally important objective of curbing the power consumption of B5G and 6G systems [8, 9, 10]. In fact, B5G and 6G systems are required to significantly improve over existing systems not only in terms of energy and spectral efficiencies, but also on data throughput, reliability, latency, coverage, and user capacity [11, 12, 13, 14], motivating the investigation of novel technologies such as millimeter-wave (mmWave)/Terahertz (THz) communications [15, 16], massive multiple-input multiple-output (mMIMO) and cell-free MIMO (CF-MIMO) [17, 18], reconfigurable intelligent surfaces (RISs) [19, 20], ISAC [3, 4, 5, 6].
@@
Among such alternatives, index modulation (IM) [21, 22, 23, 24] has the attractive and unique feature of enabling frugal utilization of resources at the transmitter, since it relies on the activation of only small subsets of the totally available resources to encode information. This characteristic also lends IM schemes great flexibility, making it transparent to other promising technologies, as demonstrated by a quickly expanding literature on the integration of IM with mMIMO, CF-MIMO, RISs and ISAC, to name a few [25, 26, 27, 28, 29, 30, 31].
@@
Generally, an IM scheme activating $P$ out of a total of $N$ resources can convey up to $\lfloor\log_{2}\binom{N}{P}\rfloor$ additional information bits on top of those conveyed over the selected $P$ resources themselves. Therefore, even with sufficiently large $N$ and an adequate $P$, very large amounts of information can be encoded in the index domain, such that IM has shown the potential to outperform fully dense ($i.e.,P=N$) systems in certain scenarios [32, 33], with the added bonus of improving the energy-efficiency and freeing the unused resources to other functionalities, such as opportunistic communications and radar/sensing [34, 35], etc.
@@
The price of the aforementioned advantages of IM are, however, the burden cast onto the receiver in terms of the computational complexity required to detect the activated resources at each given transmission instance, given all possible resource activation patterns in the codebook of size $2^{\lfloor\log_{2}\binom{N}{P}\rfloor}$. This challenge has therefore motivated work on low-complexity decoders for IM, in its many variations. To cite some examples, improvements over the brute-force maximum-likelihood (ML) search based on reduced-search methods such as lattice and sphere decoders [36, 37, 38], are able to decrease the total decoding complexity in proportion to the total codebook size, but their cost are still dependent on the combinatorial factor.
@@
More recently, methods aiming to directly reduce the order of the binomial coefficient have also been proposed. In [39], for instance, the complexity order is reduced from $\binom{N}{P}$ to $\binom{\alpha N}{P}$ with $\frac{P}{N}\leq\alpha\leq 1$, by leveraging the orthogonal matching pursuit (OMP) algorithm to operate over the probable non-zero indices of the sparse signal of the generalized quadrature spatial modulation (GQSM), which is a specific type of IM. It was shown thereby that a trade-off between computational complexity and optimal performance can be achieved by adjusting the value of $\alpha$. In [40], on the other hand, a GQSM detector leveraging the message passing (MP) algorithm with vector-valued variables under the Gaussian belief propagation (GaBP) framework is proposed, which reduces decoding cost to the order of the square root of the binomial coefficient, i.e., $\binom{N}{P}{\!\!~{}}^{\!\frac{1}{2}}$, while approaching the performance of ML solution. Finally, there exists also non-classical approaches which are independent of the binomial coefficient, including techniques based on machine learning (ML) [41, 42] and quantum computing (QC) [43]. Such alternatives have, nevertheless, their own limitations, such as the need for exhaustive off-line training or expensive and not widely-available quantum computers, respectively.
In light of all the above, we propose in this article a novel paradigm combining in which the advantages of massive IM and ISAC techniques are combined. The key idea is that in a massive multiple-input multiple-output (MIMO)-ISAC system, one can convey information by relying solely on the massively-scaled IM scheme, reducing transmit symbols to pilots, which can then be fully exploited to enable sensing functionalities. The approach implicates that such IM-enabled ISAC systems do not suffer from interference between the pilots (sensing signal) and data (information signal), thus naturally mitigating the potential performance degradation typical of ISAC schemes [44, 45]. It will be shown, in fact, that for sufficiently large sizes the proposed method outperforms ISAC systems with the same rate and power constraints designed via a “classic multiplex paradigm,” in which symbols and pilots are transmitted simultaneously, for the purpose of communication and sensing, respectively.
The only potential hurdle of the approach is, therefore, the decoding complexity at the receiver which, although prohibitive in massive set-ups under naive detection, is instead of polynomial order on other system parameters and entirely independent of the combinatorial coefficient of the codebook, thanks to the new MP detector here proposed.
While the proposed method is applicable to IM schemes of general formulation, for the sake of clarity of exposition we focus on the GQSM variation of IM, leveraging a novel unit vector decomposition (UVD) formulation of GQSM signals that enables the concurrent estimation of the index positions as independent unit vectors. Finally, several modifications to the transmitter constellation design and the detection algorithm are also provided, which enhance the performance of the detector.
Our contributions can thus be summarized as follows:
@@
  • Two enhancement modules for the proposed detector, of linear complexity, respectively based on successive interference cancellation (IC) and a conditional probability-based denoiser, are presented.
  • A new mechanism to optimize the symbol constellation for the GQSM transmitter in order to improve its UVD-based detection performance is presented.
  • In light of the above proposed methods, a new paradigm of massive MIMO systems for high-performance ISAC functionalities with controlled complexity is proposed, which can achieve better bit error rate (BER) performance than the conventional multiplexing approach, in increasingly larger system sizes under the same rate, power, and piloting constraints.
@@ II System Model
As mentioned above, many IM techniques can be found in the literature [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 46, 47, 48, 49] which address various issues of the problem such as integration with other functionalities ($e.g.$ radar and/or sensing), achievable rates, decoding complexity and, of course which resource domains to exploit and under which architecture. Without loss of generality, we consider the GQSM [46, 39] approach in particular, and leverage its system and signal models in the following derivation and analysis. It shall become clear, however, that the proposed detection scheme is applicable to many other state-of-the-art (SotA) IM schemes that can be reformulated under the UVD model presented in Section III-B.
@@ II-A GQSM System Model and Transmitter Structure
Consider a point-to-point (P2P) MIMO wireless communications system where the transmitter and the receiver are equipped with $N_{T}$ and $N_{R}$ antennas, respectively, such that the receive signal $\mathbf{y}\in\mathbb{C}^{N_{R}\times 1}$ corresponding to the transmission of the information vector $\mathbf{x}\in\mathbb{C}^{N_{T}\times 1}$ through the flat-fading MIMO channel $\mathbf{H}\in\mathbb{C}^{N_{R}\times N_{T}}$ is described by
@@
In a GQSM  [^1] Note that the classic quadrature spatial modulation (QSM) [47] is a particular case of GQSM with $P=1$. transmitter, the in-phase and quadrature (IQ) components of $P$ transmit symbols $\mathbf{s}=\{s^{\mathrm{c}}_{1},\cdots\!,s^{\mathrm{c}}_{P}\}$, taken from a complex-valued constellation $\mathcal{S}^{\mathrm{c}}$ of size $M\triangleq|\mathcal{S}^{\mathrm{c}}|$, are sparsely mapped into the respective transmit vector components $\mathbf{x}^{\mathrm{R}}\in\mathbb{C}^{N_{T}\times 1}$ and $\mathbf{x}^{\mathrm{I}}\in\mathbb{C}^{N_{T}\times 1}$ of the form
@@
$\mathbf{x}=\mathbf{x}^{\mathrm{R}}+j\mathbf{x}^{\mathrm{I}}\in\mathbb{C}^{N_{T}\times 1}.\vspace{-0.5ex}$ (2c)
@@
For the sake of clarity, consider a small system with $N_{T}=5,P=3$, such that a total of $\binom{5}{3}=10$ activation vectors are possible, namely
@@ II-B Reformulated System Model
The real domain IQ-decoupled form $\bm{y}=\bm{H}\bm{x}+\bm{w}$ presented as equation (5) is the basis of most QSM/GQSM detection algorithms, which generally seek to solve the recovery problem
@@
The problem described by equation (7) is NP-hard due to the discrete solution domain $\mathcal{X}$, which prevents classical convex optimization approaches [50], requiring instead a combinatorial search over a high-dimensional space for optimal solution. To circumvent this challenge, SotA low-complexity solutions have been proposed by leveraging various characteristics of the GQSM signal. For example, the inherent sparsity of $\bm{x}$ can be exploited by compressive sensing (CS)-based methods [51, 52, 53], and the discrete codebook structure can be utilized to design search-based methods such as sphere decoders and pruned-search decoders [54, 55, 56, 57]. Finally, a vector-valued MP method was recently proposed in [40], where the two variables $\mathbf{x}^{\mathrm{R}}$ and $\mathbf{x}^{\mathrm{I}}$ are concurrently estimated by leveraging the independent and identically distributed (i.i.d.) bivariate property of the decoupled vectors in equation (6). The bivariate vector-valued MP method of [40] was shown to reduce the search space from $R=(2^{\lfloor\log_{2}\!\binom{N_{T}}{P}\rfloor})^{2}\!\cdot\!M^{P}$ to $2\sqrt{R}=2\cdot(2^{\lfloor\log_{2}\!\binom{N_{T}}{P}\rfloor})\!\cdot\!(\!\sqrt{M\;\!})^{P}.$
@@ III Proposed UVD-based System Reformulation
In this section, the proposed unit vector decomposition (UVD)of the GQSM system model and the derivation of the novel index activation probability distributions arising from this decomposition are presented. As shall be clarified, the resultant index-wise distributions enables the design of an MP algorithm operating on a combinatorics-free signal domain.
@@ III-A Unit Vector Decomposition (UVD) of GQSM Scheme
Then, the reformulation of the transmit vector $\mathbf{x}$, the IQ-decoupled received signal vector of equation (5) becomes
@@ III-B UVD Model for Coded and Space-time Coded IM
From the equivalence among equations (17), (14) and (10), it follows once again that the UVD model introduced in (III-A) for the GQSM scheme generalizes to other IM methods in other domains and higher dimensions. For the sake of simplicity, we shall in the remainder of the article continue to use the GQSM scheme as a template of for our contribution, emphasizing however the generality of the methods as above.
@@ III-C Prior Distribution of Unit Vector Variables
IV Proposed Low-Complexity Detection for ISAC System via Massive Index Modulation
With possession of the analytical prior PMFs of the indices $k_{p}$ given in equations (18) and (23), respectively, we next design a GaBP-based MP algorithm for their estimation. In particular, we will derive the MP rules for a piloted GQSM, i.e., where $s^{\mathrm{R}}_{p},s^{\mathrm{I}}_{p}~{}\forall p$ are known, so as to enable sensing functionalities [63, 64] exploiting the symbols, as well as communication [^4] We highlight that, in principle, a bilinear MP decoder for this system could also be developed [65, 66, 64, 67], such that some of the pilots could be replaced by information-carrying symbols to increase the rate of communications, albeit at the expense of some performance degradation. Such extension is out of scope of the current article and will be pursued in a follow-up work. via the detection of activation patterns.
First, the soft-replicas for the activation vector variables $\mathbf{e}_{k^{\mathrm{R}}_{p}}$ and $\mathbf{e}_{k^{\mathrm{I}}_{p}}$ with $p\in\{1,\cdots,P\}$ are defined for the $n$-th factor node ($i.e.,n$-th pilot) with $n\in\{1,\cdots,2N_{R}\}$, as $\hat{\mathbf{e}}_{k^{\mathrm{R}}_{p}:n}$ and $\hat{\mathbf{e}}_{k^{\mathrm{I}}_{p}:n}$, respectively. The corresponding error covariance matrices of the activation vector soft-replicas are defined as

### v2

# Enabling Massive Index Modulation Systems via Combinatorics-Free Detection

@@ Abstract
Index modulation (IM) is one of the key enabling technologies for beyond fifth generation (B5G) and sixth generation (6G) wireless systems, attracting attention for its inherent energy and spectral efficiency resulting from conveying information through the indexation of the resources utilized in during signal transmission. However, a remaining critical bottleneck for large-scale IM is the consequently infeasible detection complexity of combinatoric order. Therefore in this article, in order to maximally reap the advantages of IM in large scenarios, we propose a novel message passing (MP) decoder designed under the Gaussian belief propagation (GaBP) framework exploiting a novel unit vector decomposition (UVD) of IM signals with purpose-derived novel probability distributions. The proposed method enjoys a low decoding complexity that is independent of previously prohibitive combinatorial factors, while still approaching the performance of unfeasible state-of-the-art (SotA) search-based methods. The effectiveness of the proposed approach is demonstrated via complexity analysis and numerical results for the exemplary piloted generalized quadrature spatial modulation (GQSM) systems of truly massive sizes (up to 96 antennas).
@@
Index modulation, quadrature spatial modulation, massive scale, message passing, low-complexity decoder.
@@ I Introduction
The expectations of beyond fifth generation (B5G) and sixth generation (6G) wireless systems prospect significant improvements over the existing systems in various different aspects including reliability, latency, energy and spectral efficiencies, coverage, user capacity, and more [4, 5, 6, 7], motivating the investigation of a plethora of novel technologies such as millimeter-wave (mmWave)/Terahertz (THz) communications [8, 9], massive multiple-input multiple-output (mMIMO) and cell-free MIMO (CF-MIMO) [10, 11], reconfigurable intelligent surfaces (RISs) [12, 13], and integrated sensing and communications (ISAC) [14, 15, 16, 17].
@@
Among these many novel technologies, index modulation (IM) [18, 19, 20, 21, 22] has recently gained great attention as a potential enabling technology in its inherently high throughput, energy and spectral efficiencies. Specifically, this attractive and unique feature of enabling frugal utilization of resources at the transmitter is enabled by the activation of only small subsets of the totally available resources to encode information. This characteristic also lends IM schemes great flexibility, making it transparent to other promising technologies, as demonstrated by a quickly expanding literature on the integration of IM with classic communications concepts such as orthogonal frequency division multiplex (OFDM) [26, 23, 24, 25], multiple-input multiple-output (MIMO) [27, 28, 29, 30], and multiple access [31, 32], and next-generation technologies such as RIS and ISAC [33, 34, 35, 36].
@@
Generally, an IM scheme activating $P$ out of a total of $N$ resources can convey up to $\lfloor\log_{2}\!\binom{N}{P}\rfloor$ additional information bits on top of those conveyed over the selected $P$ resources themselves, for example, via digital symbol modulation. Therefore, even with sufficiently large $N$ and an adequate $P$, significant amounts of information can be encoded in the index domain, such to highlight the potential of IM to outperform conventional systems in certain scenarios [37, 23], with the added bonus of improving the energy-efficiency and freeing the unused resources to other functionalities, such as opportunistic communications and sensing [38, 39], etc.
@@
The price of the aforementioned advantages of IM are, however, the burden cast onto the receiver in terms of the computational complexity required to detect the activated resources at each given transmission instance, given all possible resource activation patterns in the codebook of size $2^{\lfloor\log_{2}\binom{N}{P}\rfloor}$. This challenge has therefore motivated work on low-complexity decoders for IM, in its many variations. To cite some examples, improvements over the brute-force maximum-likelihood (ML) search based on reduced-search methods such as lattice and sphere decoders [40], have decreased the total decoding complexity in proportion to the total codebook size, also by exploiting the specific IM design properties [41, 24, 25]. However, such methods are still generally dependent on the combinatorial factor, prohibitive for truly massive scales.
@@
More recently, methods aiming to directly reduce the order of the binomial coefficient have also been proposed. In [42], for instance, the complexity order is reduced from $\binom{N}{P}$ to $\binom{\alpha N}{P}$ with $\frac{P}{N}\leq\alpha\leq 1$, by leveraging the orthogonal matching pursuit (OMP) algorithm to operate over the probable non-zero indices of the sparse signal of the generalized quadrature spatial modulation (GQSM), which is a specific type of IM. It was shown thereby that a trade-off between computational complexity and optimal performance can be achieved by adjusting the value of $\alpha$. On the other hand, in [43], a GQSM detector leveraging the message passing (MP) algorithm with vector-valued variables under the Gaussian belief propagation (GaBP) framework is proposed, which reduces decoding cost to the order of the square root of the binomial coefficient, i.e., $\binom{N}{P}{\!\!~{}}^{\!\frac{1}{2}}$, while approaching the performance of ML solution. Finally, there exists also non-classical approaches which are independent of the binomial coefficient, including techniques based on machine learning (ML) [44, 45] and quantum computing (QC) [46]. Such alternatives have, nevertheless, their own limitations, such as the need for exhaustive off-line training or expensive and uncommon quantum computers.
In light of all the above, we propose in this article a novel framework for IM detection which can enable a significantly reduced detection complexity that is complete independent of the prohibitive combinatorial factor inherent to IM, while still incorporating the elaborate combinatorial index codebook correlations. This is enabled by a novel unit vector decomposition (UVD) representation of IM signals, which reduce the effective signal space from the total signal vector variable into smaller unit vectors and scalars with significantly smaller effective signal spaces. Then, further aided by a novel statistical analysis of the non-trivial index-wide activation probabilities of the decoupled variables, tailored GaBP-based MP algorithms are designed to enable the concurrent estimation of the index activation information. On top of the achieved feasible complexity, the latter novel property also motivates massive IM schemes as a promising enabler of interference-free ISAC, by conducting information encoding only in the index domain whereas the resource-occupying signals can be utilized for sensing.
The effectiveness of the proposed framework is validated by simulations under the piloted GQSM variant of IM, which shows a reduced complexity where the full simulation results of unprescedented massive MIMO IM sizes (upto $96\times 96$) was feasibly obtained with even a personal computer. In addition, several modifications to the transmitter pilot constellation design and the detection algorithm are also provided, which further enhance the performance of the proposed detector.
In all, our contributions can be summarized as follows:
@@
  • Two enhancement modules for the proposed detector are presented, respectively based on successive interference cancellation (IC) and a conditional probability-based denoiser, in addition to a new mechanism to optimize the symbol constellation for the GQSM transmitter in order to improve its UVD-based detection performance.
  • Simulation results and numerical and complexity analyses of the proposed detectors of truly massive GQSM systems are provided and compared against the state-of-the-art (SotA) detectors and fully multiplexed MIMO, with sizes upto $96\times 96$ MIMO, which is unprescedented in related literature due to the infeasible detection complexity.
Notation
Complex scalars, vectors, and matrices are denoted by lowercase, boldface lowercase, and boldface uppercase characters, respectively, i.e., $x,\mathbf{x},\mathbf{X}$. ${}\Re\{{\cdot}\}$ and ${}\Im\{{\cdot}\}$ denote the real and imaginary parts of a complex variable, while the superscripts $(\cdot)^{\mathrm{R}}$ and $(\cdot)^{\mathrm{I}}$ respectively denote variables which are related to the real and imaginary parts of the system, which may still be complex-valued. Other specific notations and functions are further defined in the text as necessary.
@@ II System Model
As mentioned above, many SotA IM techniques can be found in the literature [18, 19, 20, 21, 22, 26, 23, 24, 25, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 51, 52, 53, 54, 55, 56] which address various issues of the problem such as integration with other functionalities ($e.g.$ radar and/or sensing), achievable rates, decoding complexity and, of course which resource domains to exploit and under which architecture. Without loss of generality, we exemplify the derivation and analysis of the system and signal models with the GQSM [30, 42] design [^1] It will be shown in Section III-B that the proposed design is applicable to any other type of IM-based scheme, by leveraging a UVD reformulation..
@@ II-A GQSM System Model and Transmitter Structure
Consider a point-to-point [^2] The extension of the system model and the proposed techniques to a multi-user scenario requires addressing additional non-trivial challenges, most crucially the design of multi-user IM activation codebooks and the corresponding multi-user interference-aware detector designs, and therefore will be considered in a future work. (P2P) MIMO wireless communications system where the transmitter and the receiver are equipped with $N_{T},N_{R}\!>\!1\!\in\!\mathbb{R}$ antennas, respectively, such that the receive signal $\mathbf{y}\in\mathbb{C}^{N_{R}\times 1}$ corresponding to the transmission of the information vector $\mathbf{x}\in\mathbb{C}^{N_{T}\times 1}$ through the flat-fading MIMO channel $\mathbf{H}\in\mathbb{C}^{N_{R}\times N_{T}}$ is described by
@@
In a GQSM transmitter, the in-phase and quadrature (IQ) components of $P$ transmit symbols $\mathbf{s}=\{s^{\mathrm{c}}_{1},\cdots\!,s^{\mathrm{c}}_{P}\}$, taken from a complex-valued constellation $\mathcal{S}^{\mathrm{c}}$ of size $M\triangleq|\mathcal{S}^{\mathrm{c}}|$, are sparsely mapped into the respective transmit vector components $\mathbf{x}^{\mathrm{R}}\in\mathbb{C}^{N_{T}\times 1}$ and $\mathbf{x}^{\mathrm{I}}\in\mathbb{C}^{N_{T}\times 1}$ of the form
@@
$\mathbf{x}=\mathbf{x}^{\mathrm{R}}+j\mathbf{x}^{\mathrm{I}}\in\mathbb{C}^{N_{T}\times 1}.\vspace{-1ex}$ (2c)
@@
For the sake of clarity, consider a small system with $N_{T}=5,P=3$ with $\binom{5}{3}\!=\!10$ activation vectors:
@@ II-B Reformulated System Model
The real domain IQ-decoupled form $\bm{y}=\bm{H}\bm{x}+\bm{w}$ presented as equation (5) is the basis of most quadrature spatial modulation (QSM)/GQSM detection algorithms, which generally seek to solve the recovery problem
@@
The problem described by equation (7) is NP-hard due to the discrete solution domain $\mathcal{X}$, which prevents classical convex optimization approaches [50], requiring instead a combinatorial search over a high-dimensional space for optimal solution. To circumvent this challenge, SotA low-complexity solutions have been proposed by leveraging various characteristics of the GQSM signal. For example, the inherent sparsity of $\bm{x}$ can be exploited by compressive sensing (CS)-based methods [51, 52], and the discrete codebook structure can be utilized to design search-based methods such as sphere decoders and pruned-search decoders [53, 41, 54]. Finally, a vector-valued MP method was recently proposed in [43], where the two variables $\mathbf{x}^{\mathrm{R}}$ and $\mathbf{x}^{\mathrm{I}}$ are concurrently estimated by leveraging the independent and identically distributed (i.i.d.) bivariate property of the decoupled vectors in equation (6). The bivariate vector-valued MP method of [43] was shown to reduce the search space from $R=(2^{\lfloor\log_{2}\!\binom{N_{T}}{P}\rfloor})^{2}\!\cdot\!M^{P}$ to $2\sqrt{R}=2\cdot(2^{\lfloor\log_{2}\!\binom{N_{T}}{P}\rfloor})\!\cdot\!(\!\sqrt{M\;\!})^{P}.$
@@ III Proposed UVD-based System Reformulation
In this section, the proposed unit vector decomposition (UVD)of the GQSM system model and the derivation of the novel index activation probability distributions arising from this decomposition are presented. As shall be clarified, the resultant index-wise distributions enable the design of an MP algorithm operating on a combinatorics-free signal domain.
@@ III-A Unit Vector Decomposition (UVD) of GQSM Scheme
Then, by reformulating the transmit vector $\mathbf{x}$, the IQ-decoupled received signal vector of equation (5) becomes
@@ III-B UVD Model for Coded and Space-time Coded IM
From the equivalence among equations (17), (14) and (10), it follows once again that the UVD model introduced in (III-A) for the GQSM scheme generalizes to other IM methods in other domains and higher dimensions. Therefore, we emphasize again that all of the following derivations, analyses, and proposed methods are applicable to all general types of IM. However, for the sake of simplicity, the remainder of the article will continue to use the GQSM scheme as an exemplary case.
@@ III-C Prior Distribution of Unit Vector Variables
IV Proposed Low-Complexity Detector
for Pilotted Massive Index Modulation Schemes
With possession of the analytical prior PMFs of the indices $k_{p}$ given in equations (18) and (23), respectively, we next design a GaBP-based MP algorithm for their estimation. In particular, we will derive the MP rules for a piloted GQSM, i.e., where $s^{\mathrm{R}}_{p},s^{\mathrm{I}}_{p}~{}\forall p$ are known, which can be utilized to enable secondary functionalities [^5] We highlight that, in principle, a bilinear MP decoder for this system could also be developed [65, 63, 64, 66], such that some of the pilots could be replaced by information-carrying symbols to increase the rate of communications, albeit at the expense of some performance degradation. Such extension is out of scope of the current article and will be pursued in a follow-up work. such as sensing [60, 63, 61, 62], as well as communication via the detection of activation patterns.
First, the soft-replicas for the activation vector variables $\mathbf{e}_{k^{\mathrm{R}}_{p}}$ and $\mathbf{e}_{k^{\mathrm{I}}_{p}}$ with $p\in\{1,\cdots,P\}$ are defined for the $n$-th factor node ($i.e.,n$-th pilot) with $n\in\{1,\cdots,2N_{R}\}$, as $\hat{\mathbf{e}}_{k^{\mathrm{R}}_{p}:n}$ and $\hat{\mathbf{e}}_{k^{\mathrm{I}}_{p}:n}$, respectively, as illustrated in the factor graph of Figure 1. The corresponding error covariance matrices of the activation vector soft-replicas are defined as